OpenStudy (anonymous):

solve using the elimination method -6x-6y=-12 -6x+7y=14

OpenStudy (amistre64):

you should be able to define what the elimination method does .....

OpenStudy (anonymous):

change the bottom equation to 6x - 7y = -14 add both equations together and simply. you get -13y = -26 y = 2 plug in 2 for y. -6x - 6(2) = -12 -6x =0 x = 0